a body of mass 1 kg is allowed to fall free under gravity. find the kinetic energy of the body 5s after its starts falling

Grade:9

2 Answers

Ajay Mandal
7 years ago
Initial velocity = 0Acceleration = 9.8 m/s^2Find final velocity after 5 second.Then kinetic energy = 1/2 mv^2
sehwag
7 years ago
V=u+at. V=0+9.8×5 =49. K,e= mv2/2= 2401/2=1200.5 first we find v then we substitute the value of v in equation u =0 because it is initial velocity